The Cost of Equity and Flotation Costs Messman Manufacturing will issue common stock to the public for $40. The expected dividend and the growth in dividends are $3.25 per share and 4%, respectively. If the flotation cost is 15% of the issue's gross proceeds, what is the cost of external equity, re? Round your answer to two decimal places. %
Answer #1

Price of stock net of flotation cost =40*(1-15%) =34

Cost of external equity:-

=Expected dividend/price+growth

=3.25/34+4%

=13.56%